OpenStudy (anonymous):

Jackson is on a new exercise regimen. To begin, his trainer does not want him to complete more than a total of 27 lunges and sit-ups. For every lunge that he completes, Jacob burns 30 calories. He burns an additional 20 calories for every sit-up. His goal is to burn at least 620 calories during a workout session. Which is a combination of lunges and sit-ups that Jacob can complete in a session? 7 lunges and 19 sit-ups 11 lunges and 15 sit-ups 10 lunges and 14 sit-ups 12 lunges and 19 sit-ups

OpenStudy (anonymous):

Cool thanks, I will try and figure out some if I can with that equation. Thanks.

OpenStudy (anonymous):

So just use the process of elimination? I tried doing that but none of my math turned out to be any answer choice given :/

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

whats so funny dude?

OpenStudy (anonymous):

hmm equations

OpenStudy (anonymous):

let me explain

OpenStudy (anonymous):

so the answer is b?

OpenStudy (anonymous):

Go ahead you can probably help her way more that I can.

OpenStudy (anonymous):

let lunges = x and sit ups = y so his trainer has said not to do total more than 27 so x+y =27 and he should burn atleast 620 calories so 30 calories for lunges and 20 for situps so you have 30x+20y =620 solve these 2 equations you will get answer

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Goodness...I knew how to do that and totally forgot about the other equation :P Lol I always do stuff the hard way cause i forget the easier way.

OpenStudy (anonymous):

i am not so good at substitution but I am leaning towards answer choice b. am i correct?

OpenStudy (anonymous):

star hunterrr what will you do if there are no options given

OpenStudy (anonymous):

lol ^

OpenStudy (anonymous):

Well I work backwards a lot, so I guess I will fail. I know I need to practice more. but for now, I just wanna get this review done :) and make sure I am ready for my exam tomorrow.

OpenStudy (anonymous):

well you cant solve the equations then are you?

OpenStudy (anonymous):

okay do this x+y =27 x= 27-y put value of x in 2nd equation and find y

OpenStudy (anonymous):

Well now i am doubting myself. I worked through the steps and 630 is what i got. So. am i right? about the answer being b?

OpenStudy (anonymous):

11+25 = 26

OpenStudy (anonymous):

not 27

OpenStudy (anonymous):

well none of them totals 27 :|

OpenStudy (anonymous):

yeah that's where i am getting stuck. I keep plugging in all the answer choices but none of them seem to make sense.

OpenStudy (anonymous):

i guess it is right not more than 27 and atleast 620 so you can go for B

OpenStudy (anonymous):

that was my rationale